OpenStudy (anonymous):

Consider the line that passes through the point (3, -6) and has a slope of 4. Part 1: Write the equation of this line using point-slope form. (2 points) Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. (2 points) Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work. (2 points)

OpenStudy (anonymous):

Part 1: y1= -6, x1= 3 and m4 y-1 (-6) =4 (x-3) y+6 = 4 (x-3) Part 2: y+6 = 4(x-3) y+6 = 4x - 12 y+4x - 18 Part 3: Y=4x -18 -4x + y = -18 (4x+y=-18) x-1 4x - y =18 Is this right?